Check my workCheck My Work button is now enabled

Item 10

Item 10 1.4 points Item Skipped

Exercise 5-17 Break-Even and Target Profit Analysis [LO5-4, LO5-5, LO5-6]

Outback Outfitters sells recreational equipment. One of the company’s products, a small camp stove, sells for $50 per unit. Variable expenses are $32 per stove, and fixed expenses associated with the stove total $108,000 per month.

Required:

1. What is the break-even point in unit sales and in dollar sales?

2. If the variable expenses per stove increase as a percentage of the selling price, will it result in a higher or a lower break-even point? (Assume that the fixed expenses remain unchanged.)

3. At present, the company is selling 8,000 stoves per month. The sales manager is convinced that a 10% reduction in the selling price would result in a 25% increase in monthly sales of stoves. Prepare two contribution format income statements, one under present operating conditions, and one as operations would appear after the proposed changes.

4. Refer to the data in (3) above. How many stoves would have to be sold at the new selling price to attain a target profit of $35,000 per month?
Answer #1

1) Break even unit = 108000/(50-32) = 6000 Units

Break even sales = 6000*50 = $300000

2) Like selling price increase 10% = 50*1.1 = 55

Variable cost = 32*1.1 = 35.20

Break even point = 108000/(55-35.20) = 5455

So its lower break even point

3) Contribution margin income statement

Present Proposed
Total Per unit Total Per unit
Sales 400000 50 450000 45
Variable cost 256000 32 320000 32
Contribution margin 144000 18 130000 13
Fixed cost 108000 108000
Net income 36000 22000

4) Required unit = (108000+35000)/13 = 11000 Units